# Fast and Accurate Collimator-Detector Response Compensation in High-Energy SPECT Imaging with 1D Convolutions and Rotations

### v1

@@ Abstract
Objective: Modeling of the collimator-detector response (CDR) in SPECT reconstruction enables improved resolution and more accurate quantitation, especially for higher energy imaging (e.g. 177Lu and 225Ac). Such modeling, however, can pose a significant computational bottleneck when there are substantial components of septal penetration and scatter in the acquired data, since a direct convolution-based approach requires large 2D kernels. The present work presents an alternative method for fast and accurate CDR compensation using a linear operator built from 1D convolutions and rotations (1D-R). To enable open-source development and use of these models in image reconstruction, we release a SPECTPSFToolbox repository for the PyTomography project on GitHub. Approach: A 1D-R CDR model was formulated, and subsequently fit to Monte Carlo 440 keV point source data representative of emissions in 225Ac imaging. Computation times of (i) the proposed 1D-R model and (ii) a traditional model that uses 2 dimensional convolutions (2D) were compared for typical SPECT matrix sizes. Both CDR modeling techniques were then used to reconstruct 225Ac phantom and patient data, and were compared by quantifying total counts in hot regions of interest (ROIs) and activity contrast between hot ROIs and background regions. Results: The 1D-R and 2D CDR models were created using the SPECTPSFToolbox. For typical matrix sizes in SPECT reconstruction, application of the 1D-R model provides a two-fold computational speed-up over the 2D model. Only small differences between the 1D-R and 2D models (order of $1\%$ ) were obtained for count and contrast quantification in select ROIs. Significance: A technique for CDR modeling in SPECT was proposed that (i) significantly speeds up reconstruction times, and (ii) yields nearly identical reconstructions to traditional 2D convolution based CDR techniques. The released toolbox will permit open-source development of similar models for different isotopes and collimators.
@@ 1 Introduction
Single photon emission computed tomography (SPECT) imaging is an in vivo modality of significant value in different clinical applications [1, 2, 3]. In particular, quantitative SPECT imaging holds great potential and value in the field of theranostics, wherein image quantification and dosimetry can be performed to improve radiopharmaceutical therapies (RPTs). Routine SPECT-based dosimetry could unveil the relationships between tumor-response and healthy organ complications with absorbed dose, therefore enabling personalized treatments that maximize tumor dose while minimizing toxicity to organs at risk [4, 5]. As a recent and well-known example, the success of 177Lu-PSMA-617 based RPTs in the VISION [6] and TheraP [7] randomized control trials has lead to expanded research and clinical translation efforts for wide-scale deployment, as well as pursuit of absorbed dose estimation as a tool towards personalized treatments [8, 9].
@@
SPECT-based dosimetry relies on accurate quantification of activity distribution from SPECT images, which is strongly dependent on the system model used in image reconstruction. The system model is a linear operator that predicts the expectation of the acquired data given a 3D isotope distribution, accounting for phenomena such as photon attenuation in the patient. An important aspect of the model is collimator detector response (CDR) modeling, which estimates image blurring caused by the collimator and the scintillation crystals [10, 11, 12]. Computation of the CDR is typically the most computationally expensive operation in image reconstruction.
@@
The CDR can be characterized by the detector response from a point source of activity; this is denoted the point spread function (PSF). The PSF can be decomposed into multiple components. The intrinsic response function (IRF) characterizes the inability of the scintillation crystals to precisely localize the point of interaction, and is sufficiently modeled using a Gaussian function. The collimator response for parallel hole collimators results from the inability of the collimator to accept only photons travelling perpendicular to the detector; it consists of three components: (i) the geometric response function (GRF) [11, 13, 14, 15], which describes photons that travel through the collimator holes without penetrating or interacting with the septa, (ii) the septal penetration response function (SPRF), which describes the contribution from photons that travel through the collimator without being attenuated, and (iii) the septal scatter response function (SSRF), which consists of photons that interacted and scattered within the collimator and were subsequently detected in the scintillator.
@@
Selection of collimator parameters is an important aspect of SPECT imaging [16]. As the collimator becomes thicker and the diameter of the collimator bores becomes narrower, the relative contribution from the SPRF and SSRF decreases due to the increased attenuation probability for photons not travelling perpendicular to the detector. In this case, (i) the detector resolution improves and (ii) the net point spread function (PSF) is dominated by the IRF + GRF and can be reasonably approximated using a 2D Gaussian function. A trade-off of having a collimator with high septal thickness and small hole diameters, however, is a decrease in detector sensitivity; the corresponding implications for quantitative imaging is a decrease in precision or longer patient scan times. This trade-off must be independently considered for each isotope. For 177Lu labeled radiopharmaceuticals, where the imaged photons are 208208208~{}208keV, a commercially labeled “medium energy” collimator configuration (i) yields a reasonable count rate and (ii) adequately minimizes the SPRF and SSRF components. In this situation, a 2D Gaussian PSF model is sufficient for image reconstruction [17, 18].
@@
Many commercially available reconstruction software only offer Gaussian PSF modeling, and thus implicitly assume the SPRF and SSRF are negligible. There are multiple advantages to this. Firstly, the computational advantage of 2D Gaussian PSF modeling is that 2D Gaussian convolution is separable into use of two perpendicular 1D Gaussian convolutions, which are less computationally expensive to implement. Secondly, the Gaussian PSF model used to model the GRF+IRF can be obtained analytically for any photon emission energy and standard collimator shapes, and thus does not require lookup tables. However, when the PSF has significant contributions from the SPRF and SSRF, 2D Gaussian PSF modeling fails to capture all the features of the PSF. This is typically an issue with radioisotopes that emit high photon energies, such as $\alpha$-emitters like 225Ac.
@@
Use of $\alpha$-emitters in radiopharmaceutical therapies presents a major and exciting frontier, due to the high linear energy transfer (LET) associated with $\alpha$ emissions [19, 20, 21, 22, 23, 24]. In preclinical studies, simultaneous treatment with 177Lu-PSMA-617 and 225Ac-PSMA-617 compared to 177Lu-PSMA-617 alone resulted in significantly reduced tumor growth [25]. Kratochwil et al. [26] applied 225Ac-PSMA-617 to patients with metastatic castration resistant prostate cancer (mCRPC) who previously exhausted 177Lu-PSMA-617 treatment. Certain patients achieved a full response, with prostate specific antigen (PSA) levels in one patient decreasing from 419 ng/mL to below 0.1 ng/mL. At the time of writing, there are ongoing studies of 225Ac based radiopharmaceuticals looking at dose escalation [27, 28], fractionation [29], and safety and efficacy [30]. A recent meta analysis [31] elaborates further on these clinical trials, discussing toxicities and other challenges with these treatments. Quite recently, targeted $\alpha$ therapy with 213Bi has been shown to reduce amyloid plaque concentration in male mice [32], eluding to a potential treatment option for Alzheimer disease.
@@
The reduction of reconstruction times in medical imaging remains an important research topic. Tsai et al. [35] recently showed that the limited-memory Broyden-Feltcher-Goldfarb-Shannon algorithm with box constraints and a diagonal preconditioner (L-BFGS-B-PC) was able to converge several times faster than the one step late expectation maximum (OSL-EM) algorithm in Positron Emission Tomography (PET) reconstruction because less projection operations were required. As another example, in Chun et al. [34], long reconstruction times for PSF modeling of 131I were partially remedied by using fast fourier transform (FFT) based 2D convolutions. By contrast, in the present work, we seek to reduce reconstruction times by using a PSF model that incorporates (i) 1D convolutions and (ii) rotations. It will be shown that this model significantly reduces computational time compared to standard and FFT based 2D convolutions, and consequently significantly reduces the time required for SPECT image reconstruction of 225Ac. The method is evaluated on phantom data and patient data, and we show that the proposed model yields identical results to Monte Carlo based 2D convolution techniques. All models considered in this paper are implemented using the GPU-accelerated functionality of PyTorch.
@@ 2 Materials and Methods
In Section 2.1, the mathematical formalism of the PSF model is outlined, and the reconstruction protocol for all acquired data in subsequent sections is established. Section 2.2 describes how the model is fit to SIMIND [37] Monte Carlo 225Ac point source data acquisitions simulated at various source-detector distances. The computational time of the model is benchmarked and compared to 2D PSF modeling and Gaussian PSF modeling. In Section 2.3, the model is then used for reconstruction of (i) an 225Ac phantom consisting of spherical inserts and (ii) a patient receiving 225Ac-PSMA-617 treatment for metastatic prostate cancer. All computation was performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a Tesla V100 GPU.
@@ 2.1.1 PSF Modeling
$K^{(y;d;b)}f\equiv\sum_{y^{\prime}}f(x,y^{\prime};d)k(y-y^{\prime};d)$ (2)
@@
where $(x^{\prime},y^{\prime})$ is the source position on a plane parallel to the detector in 3D space, $\mathrm{psf}(...,d)$ is a kernel that yields the point spread function (PSF) at a distance $d$ from the detector, and $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)$ is the probability that photons traveling from $(x,y,d)$ to detector coordinate $(x^{\prime},y^{\prime})$ are not attenuated. Under the assumption that the attenuation probabilities vary little across the PSF, it follows that $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)\approx p_{\mathrm{att}}(x,y,x,y,d)$ and it can simply be expressed as $p_{\mathrm{att}}(x,y,d)$. Defining $f^{\prime}\equiv p_{\mathrm{att}}f$ as the “attenuation-adjusted” image, Equation 4 can then be rewritten in operator form (Equation 3) as
@@
Since this convolution operation is typically the bottleneck of SPECT system matrix modeling and image reconstruction, it is of interest to look for techniques to reduce the computation time. Under conditions of no septal penetration and scatter, the CDR is dominated by the GRF: $K_{\mathrm{PSF}}^{(x,y;d;b)}$ can then be sufficiently approximated using a 2D Gaussian convolution $K_{G}^{(x,y;d;b)}$ where the kernel is given by
@@
Convolution with a Gaussian function has computational advantages since it can be decomposed into successive application of two perpendicular 1D kernels via $K_{G}^{(x,y;d;b)}=K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$; 1D convolution is significantly more computationally efficient than 2D. Unfortunately, the decomposition of a 2D kernel into two 1D kernels is not mathematically possible when the CDR contains significant SPRF and SSRF components. 2D convolution, however, is not the only way to implement Equation 5. Owing to the discrete rotational symmetries and features of the anisotropic PSFs obtained with typical SPECT collimators, we propose the following “1D rotation” model (abbreviated as “1D-R”) for the PSF operator:
$M_{\mathrm{1D-R}}^{(x,y;d;b)}\equiv\left(\sum_{\theta\in\Theta_{t}}\mathcal{R}^{-1}_{\theta}K^{(x;d;b)}_{t}\mathcal{R}_{\theta}+K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}+1\right)K^{(x;d;b)}_{G}K^{(y;d;b)}_{G}$ (8)
where $\mathcal{R}_{\theta}$ is a rotation operator that implements rotation about an axis in the $d$ direction by angle $\theta$, and $\Theta_{t}$ correspond to the angle of the septal penetration tails (equal to $\{0,\pi/3,2\pi/3\}$ for 225Ac). The derived linear operator only makes use of 1D convolutions and rotations, since they often require less computational time on GPU compared to 2D convolutions. The kernels, modeling the components of the CDR (as shown in Figure 1), are selected as follows:
@@
    This kernel is used to build geometric component of the PSF.
@@
    $A_{t}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (13)
    $\sigma_{t}(d,b)=1+b_{4}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{5}^{2}}-|b_{5}|\right)$ (14)
    where $d_{\mathrm{min}}$ is the source-detector distance used in the PSF fit. $f_{t}(x)$ is a discrete array of numbers that is linearly interpolated between its fixed points; these points are also considered to be part of the hyperparameters $b$. This kernel is used to build the septal penetration component of the PSF.
@@
    $A_{B}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (16)
    $\sigma_{B}(d,b)=1+b_{4}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{5}^{2}}-|b_{5}|\right)$ (17)
@@
    This kernel is used to build the septal scatter component of the PSF.
Figure 1: 440 keV PSF obtained by simulating a point source at a source-detector distance of 30 cm using SIMIND. Labeled scatter components that refer to the different terms of Equation 8 are shown.
@@
In subsequent sections, the proposed 1D-R model is evaluated against a Monte Carlo 2D model where kernel data is obtained using SIMIND [37]; point source projection data is normalized at a large number of source-detector distances. The corresponding PSF operator, abbreviated as “2D”, is defined as
$M_{\mathrm{2D}}^{(x,y;d;b)}\equiv K_{\mathrm{true}}^{x,y;d}$ (19)
where the corresponding kernel $k_{\mathrm{true}}^{x,y;d}$ corresponds to normalized projection data at each source-detector distance $d$.
The publicly SPECTPSFToolbox repository is structured to facilitate the customizability of these models. For the 1D-R model, components are implemented via separate class instances that are subsequently added and multiplied together to obtain the final form of Equation 8. The functional form and hyperparameters of the amplitude/scaling, such as that of Equation 10, are left arbitrary for the user. The 2D model (Equation 19) is implemented via a class instance that receives a stack of PSF kernels with associated source-detector distances: when used in image reconstruction, the model automatically chooses the PSF kernel closest to each source-detector distance in the reconstruction problem.

### v2

# Computationally Efficient Collimator-Detector Response Compensation in High Energy SPECT using 1D Convolutions and Rotations

@@ Abstract
Objective: Modeling of the collimator-detector response (CDR) in SPECT reconstruction enables improved resolution and accuracy, and is thus important for quantitative imaging applications such as dosimetry. The implementation of CDR modeling, however, can become a computational bottleneck when there are substantial components of septal penetration and scatter in the acquired data, since a direct convolution-based approach requires large 2D kernels. This work proposes a 1D convolution and rotation-based CDR model that reduces reconstruction times but maintains consistency with models that employ 2D convolutions. To enable open-source development and use of these models in image reconstruction, we release a SPECTPSFToolbox repository for the PyTomography project on GitHub. Approach: A 1D/rotation-based CDR model was formulated and subsequently fit to Monte Carlo point source data representative of 177Lu, 131I, and 225Ac imaging. Computation times of (i) the proposed 1D/rotation-based model and (ii) a traditional model that uses 2D convolutions were compared for typical SPECT matrix sizes. Both CDR models were then used in the reconstruction of Monte Carlo, physical phantom, and patient data; the models were compared by quantifying total counts in hot regions of interest (ROIs) and activity contrast between hot ROIs and background regions. Results: For typical matrix sizes in SPECT reconstruction, application of the 1D/rotation-based model provides a two-fold computational speed-up over the 2D model when running on GPU. Only small differences between the 1D/rotation-based and 2D models (order of $1\%$) were obtained for count and contrast quantification in select ROIs. Significance: A technique for CDR modeling in SPECT was proposed that (i) significantly speeds up reconstruction times, and (ii) yields nearly identical reconstructions to traditional 2D convolution based CDR techniques. The released toolbox will permit open-source development of similar models for different isotopes and collimators.
@@ 1 Introduction
Single photon emission computed tomography (SPECT) is an in vivo imaging modality, possessing value for various clinical applications [1, 2, 3]. In particular, quantitative SPECT imaging has applications in the field of theranostics, wherein image quantification and dosimetry can be performed to improve radiopharmaceutical therapies (RPTs). Routine SPECT-based dosimetry could unveil the relationships between tumor-response and healthy organ complications with absorbed dose, therefore enabling personalized treatments that maximize tumor dose while minimizing toxicity to organs at risk [4, 5]. As a recent and well-known example, the success of 177Lu-PSMA-617 based RPTs in the VISION [6] and TheraP [7] randomized control trials has lead to expanded research and clinical translation efforts for wide-scale deployment, as well as pursuit of absorbed dose estimation as a tool towards personalized treatments [8, 9].
@@
SPECT-based dosimetry relies on accurate quantification of activity distribution from SPECT images, which is strongly dependent on the system model used in image reconstruction. The system model is a linear operator that predicts the expectation of the acquired data given a 3D isotope distribution, accounting for phenomena such as photon attenuation in the patient. An important aspect of the model is collimator detector response (CDR) modeling, which estimates image blurring caused by the collimator and the detector [10, 11, 12]. Computation of the CDR is typically the most computationally expensive operation in image reconstruction.
@@
The CDR can be characterized by the system’s response to a point source of activity; this is denoted the point spread function (PSF). The PSF can be decomposed into multiple components. The intrinsic response function (IRF) characterizes the uncertainty on the point of interaction within the detecting material (e.g. an NaI scintillator), and is sufficiently modeled using a Gaussian function. The collimator response for parallel hole collimators results from the inability of the collimator to accept only photons travelling perpendicular to the detector; it consists of three components: (i) the geometric response function (GRF) [11, 13, 14, 15], which describes photons that travel through the collimator holes without penetrating or interacting with the septa, (ii) the septal penetration response function (SPRF), which describes the contribution from photons that travel through the collimator without being attenuated, and (iii) the septal scatter response function (SSRF), which consists of photons that interacted and scattered within the collimator and were subsequently detected.
@@
The selection of collimator parameters is an important aspect of SPECT imaging [16]. A collimator with longer, narrower bores, and thicker septal thickness reduces the contribution from the SPRF and SSRF by increasing the probability of attenuation for photons not traveling perpendicular to the detector. If the collimator geometry is optimized as such, the net point spread function (PSF) is dominated by the IRF + GRF and can be reasonably approximated using a 2D Gaussian function. A trade-off of having a thicker collimator with smaller hole diameters, however, is a decrease in detector sensitivity; the corresponding implications for quantitative imaging is a decrease in precision or longer patient scan times. This trade-off must be independently considered for each isotope. For 177Lu labeled radiopharmaceuticals, where the 208208208~{}208keV photopeak is typically used for imaging, a commercially labeled “medium energy” collimator configuration (i) yields a reasonable count rate and (ii) adequately minimizes the SPRF and SSRF components. In this situation, a 2D Gaussian PSF model is sufficient for image reconstruction [17, 18].
@@
To the knowledge of the authors, all presently available commercial reconstruction software offer Gaussian PSF modeling, and thus implicitly assume the SPRF and SSRF are negligible. There are multiple advantages to this. Firstly, the computational advantage of 2D Gaussian PSF modeling is that 2D Gaussian convolution is separable into use of two perpendicular 1D Gaussian convolutions, which are less computationally expensive to implement. Secondly, the Gaussian PSF model used to model the GRF+IRF can be obtained analytically for any photon emission energy and standard collimator shapes, and thus does not require lookup tables. However, when the PSF has significant contributions from the SPRF and SSRF, 2D Gaussian PSF modeling fails to capture all the features of the PSF. This is typically an issue with radioisotopes that emit high photon energies, such as $\alpha$-emitters like 225Ac.
@@
Use of $\alpha$-emitters in radiopharmaceutical therapies presents an exciting frontier due to the high linear energy transfer (LET) associated with $\alpha$ particles, which may yield benefits for tumor eradication [19, 20, 21, 22, 23, 24]. In preclinical studies, simultaneous treatment with 177Lu-PSMA-617 and 225Ac-PSMA-617 compared to 177Lu-PSMA-617 alone resulted in significantly reduced tumor growth [25]. Kratochwil et al. [26] applied 225Ac-PSMA-617 to patients with metastatic castration resistant prostate cancer (mCRPC) who previously exhausted 177Lu-PSMA-617 treatment. Certain patients achieved a full response, with prostate specific antigen (PSA) levels in one patient decreasing from 419 ng/mL to below 0.1 ng/mL. At the time of writing, there are ongoing studies of 225Ac based radiopharmaceuticals looking at dose escalation [27, 28], fractionation [29], and safety and efficacy [30]. A recent meta analysis [31] elaborates further on these clinical trials, discussing toxicities and other challenges with these treatments. Quite recently, targeted $\alpha$ therapy with 213Bi has been shown to reduce amyloid plaque concentration in male mice [32], eluding to a potential treatment option for Alzheimer disease.
@@
The reduction of reconstruction times in medical imaging remains an important research topic. Tsai et al. [35] recently showed that the limited-memory Broyden-Feltcher-Goldfarb-Shannon algorithm with box constraints and a diagonal preconditioner (L-BFGS-B-PC) was able to converge several times faster than the one step late expectation maximum (OSL-EM) algorithm in Positron Emission Tomography (PET) reconstruction because less projection operations were required. As another example, Chun et al. [34] reduced reconstruction times of 131I by using fast fourier transform (FFT) based 2D convolutions for PSF modeling. By contrast, in the present work, we seek to reduce reconstruction times by using a PSF model that incorporates (i) 1D convolutions and (ii) rotations. It will be shown that this model runs faster than standard and FFT based 2D convolutions when implemented on a GPU, and can thus be used to speed up image reconstruction times. The method is evaluated on Monte Carlo phantom data of 177Lu, 131I, and 225Ac, as well as physical phantom and patient data of 225Ac. In each case, the method demonstrates near-identical recovery coefficients and contrast in regions of interest compared to MC-based 2D convolution techniques.
@@ 2 Materials and Methods
Section 2.1 outlines the mathematical formalism of the PSF model and establishes the reconstruction protocol for all acquired data presented in subsequent sections. Section 2.2 describes how the model is fit to Monte Carlo (MC) 177Lu, 131I, and 225Ac point source data acquisitions simulated at various source-detector distances. The computational time of the model is benchmarked and compared to 2D PSF modeling and Gaussian PSF modeling. In Section 2.3, the model is then used for reconstruction of (i) MC simulated 177Lu, 131I, and 225Ac phantoms, (ii) a physical 225Ac phantom, and (iii) a patient receiving 225Ac-PSMA-617 treatment for metastatic prostate cancer. All computation was performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a Tesla V100 GPU.
All MC data was generated using SIMIND [37]. In order to capture the effects of septal penetration and scatter, SIMIND has an optional routine based on the delta-scattering technique [38], which has been validated for high-energy isotopes such as 131I [39]. This SIMIND routine has been used in previous studies to generate stacks of PSFs at various source-detector distances for MC-based 225Ac PSF modeling in custom reconstruction [23], since the reconstruction software of the scanner manufacturers does not provide comprehensive PSF modeling of high energy isotopes.
@@ 2.1.1 PSF Modeling
$K^{(y;d;b)}f\equiv\sum_{y^{\prime}}f(x,y^{\prime};d)k(y-y^{\prime};d;b)$ (2)
@@
where $(x^{\prime},y^{\prime})$ is the source position on a plane parallel to the detector in 3D space, $\mathrm{psf}(...,d)$ is a kernel that yields the point spread function (PSF) at a distance $d$ from the detector, and $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)$ is the probability that photons traveling from $(x,y,d)$ to detector coordinate $(x^{\prime},y^{\prime})$ are not attenuated. If the attenuation probabilities are approximately constant for different photon impingement angles permitted on the detector (e.g. valid when the PSF is small), it follows that $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)$ can be replaced with $p_{\mathrm{att}}(x,y,x,y,d)$, which can simply be written as $p_{\mathrm{att}}(x,y,d)$. Under this assumption, $f^{\prime}\equiv p_{\mathrm{att}}f$ can be defined as the “attenuation-adjusted” image, and Equation 4 can then be rewritten in operator form (Equation 3) as
@@
Since this convolution operation often forms the bottleneck of SPECT system matrix modeling and image reconstruction, it is of interest to look for techniques to reduce the computation time. Under conditions of no septal penetration and scatter, the CDR is dominated by the GRF: $K_{\mathrm{PSF}}^{(x,y;d;b)}$ can then be sufficiently approximated using a 2D Gaussian convolution $K_{\mathrm{2DG}}^{(x,y;d;b)}$ where the kernel is given by
@@
Convolution with a Gaussian function has computational advantages since it can be decomposed into successive application of two perpendicular 1D kernels via $K_{\mathrm{2DG}}^{(x,y;d;b)}=K_{\mathrm{1DG}}^{(x;d;b)}K_{\mathrm{1DG}}^{(y;d;b)}$; 1D convolution is significantly more computationally efficient than 2D. Unfortunately, the decomposition of a 2D kernel into two 1D kernels is not mathematically possible when the CDR contains significant SPRF and SSRF components. 2D convolution, however, is not the only way to implement Equation 5. Owing to the discrete rotational symmetries and features of the anisotropic PSFs obtained with parallel hole SPECT collimators, we propose the following “1D rotation” model (abbreviated as “1DR”) for the PSF operator:
$M_{\mathrm{1DR}}^{(x,y;d;b)}\equiv\left(\sum_{\theta\in\Theta_{t}}\mathcal{R}^{-1}_{\theta}K^{(x;d;b)}_{t}\mathcal{R}_{\theta}+K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}+1\right)K^{(x;d;b)}_{G}K^{(y;d;b)}_{G}$ (8)
where $\mathcal{R}_{\theta}$ is a rotation operator that implements rotation about an axis in the $d$ direction by angle $\theta$, and $\Theta_{t}$ correspond to the angle of the septal penetration tails (equal to $\{0,\pi/3,2\pi/3\}$ for hexagonal collimators). This derived linear operator only makes use of 1D convolutions and rotations, since they require less computational time on GPU compared to 2D convolutions. The kernels, modeling the components of the CDR (as shown in Figure 1), are selected as follows:
@@
    This kernel is used to build geometric component of the PSF. The terms $A_{G}(d,b)$ and $\sigma_{G}(d,b)$ are able to capture the decaying amplitude and spreading out of the PSF at larger distances $d$.
@@
    $A_{t}(d,b)=b_{7}e^{-b_{8}d}+b_{9}e^{-b_{10}d}$ (13)
    $\sigma_{t}(d,b)=1+b_{11}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{12}^{2}}-|b_{12}|\right)$ (14)
    where $d_{\mathrm{min}}$ is the source-detector distance used in the PSF fit. $f_{t}(x)$ is a discrete array of numbers that is linearly interpolated between its fixed points. This kernel models the tails of the septal penetration PSF component. Similar to the geometric component, the terms $A_{t}(d,b)$ and $\sigma_{t}(d,b)$ are such that the amplitude decays and the component spreads out at larger distances $d$. The function $f_{t}(x/\sigma_{t}(d,b))$ is initialized as a mono-exponential function to match the decay of the tails observed in Figure 2.
@@
    $A_{B}(d,b)=b_{13}e^{-b_{14}d}+b_{15}e^{-b_{16}d}$ (16)
    $\sigma_{B}(d,b)=1+b_{17}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{18}^{2}}-|b_{18}|\right)$ (17)
@@
    This kernel is used to build the septal scatter component of the PSF. Similar to the previous two components, the terms $A_{B}(d,b)$ and $\sigma_{B}(d,b)$ capture the amplitude decay and spreading out of the component at larger distances $d$. The function form of $f_{B}(x)$ was selected based on the approximate decay of septal scatter seen in Figure 2.
Figure 1: (a) Top view of collimator, where $w_{b}$ is the collimator hole width. (b) Side view of collimator, where $L_{b}$ is the length of the collimator. Three different photon detection paths are shown as black arrows; from left to right they correspond to septal scatter, septal penetration, and geometric component respectively. (c) 440 keV PSF obtained by simulating a point source at a source-detector distance of 30 cm using SIMIND. Labeled scatter components that refer to the different terms of Equation 8 are shown.
@@
PSF modeling can also be accomplished using a generic 2D kernel; the model can be expressed as
$M_{\mathrm{2D}}^{(x,y;d;b)}\equiv K_{\mathrm{2D}}^{(x,y;d)}$ (19)
where the corresponding kernel $k_{\mathrm{2D}}^{(x,y;d)}$ might be obtained from real data measurements, MC measurements, or an analytical model based on measurements (such as in [35]). In subsequent sections, the following three methods are compared:
  1. 1.
    The 2D Gaussian model of Equation 6, with parameters analytically obtained from collimator dimensions. Denoted “2DG”, this is the technique used by available commercial reconstruction software.
  2. 2.
    A generic 2D Monte Carlo model (Equation 19) obtained directly from a stack of normalized point source MC data at various source-detector distances. Denoted “2D-MC”, this is the same technique used by Delker et. al. [23] for reconstruction of 225Ac patient data; it will be used as a standard against which method (iii) is compared.
  3. 3.
    The proposed 1DR model (Equation 8) with parameters obtained via fitting to MC point source data, and denoted “1DR-MC”. Since it is fit to the data used for method (ii), an ideal model would yield identical reconstructions to those obtained with method (ii).
It should be emphasized that method (i) does not require any empirical data, while methods (ii) and (iii) use MC point source data (hence the inclusion of “-MC” in the acronyms). Methods (ii) and (iii), however, can also be implemented using real point source measurements if that data is available.
The publicly available SPECTPSFToolbox repository is structured to facilitate the customizability of these models. For the 1DR-MC model, components are implemented via separate class instances that are subsequently added and multiplied together to obtain the final form of Equation 8. The functional form and hyperparameters of the amplitude/scaling, (e.g. Equation 10) are left arbitrary for the user; this permits full customization of PSF models. The 2D-MC model is implemented via a class instance that receives a stack of PSF kernels with associated source-detector distances: when used in image reconstruction, the model automatically chooses the PSF kernel closest to each source-detector distance in the reconstruction problem.